# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 9.1 Regularizing the graph and structural properties
We say that $G$ is in 5-base position (or 5BP for short) if the following conditions hold:
@@
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
@@
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Note that, if $G$ is in 5BP, then any 5-vertex $u$ only has neighbors of degree 4 and 5 (else we would have $R(u)\geq 1$). Unlike in the degree-3 case, we cannot ensure that $G$ has no four-cycles. However, we can show that the four-cycles are heavily restricted.
@@
If $G$ is in 5BP, then any pair of non-adjacent vertices $x,y$ have $\operatorname{codeg}(x,y)\leq 2$.
@@
Suppose $\deg(x)\geq\deg(y)$ and $\operatorname{codeg}(x,y)\geq 3$. If $\deg(x)=3$, then $\{x,y\}$ would be an indset with surplus one, contradicting that $G$ is simplified. If $\deg(y)=4$, then $\operatorname{shad}(N[y])\leq 0$ (due to vertex $x$), contradicting that $G$ is in 5BP. So we suppose $\deg(x)=\deg(y)=5$. Again, if $\operatorname{codeg}(x,y)\geq 4$ then $\operatorname{shad}(N[y])\leq 0$ due to vertex $x$, so we suppose that $\operatorname{codeg}(x,y)=3$ exactly. Then $R(N(x)\cap N(y))\geq 2$ due to non-adjacent 2-vertices $x,y$. So $R(x,y)+R(N(x)\cap N(y))+\operatorname{codeg}(x,y)\geq 5$ contradicting that $G$ is in 5BP. ∎
@@
If $G$ is in 5BP, then any 4-cycle $x_{1},x_{2},x_{3},x_{4}$ satisfies the following properties:
@@
For the first result, suppose that $\deg(x_{1})=3$. If $x_{1}\sim x_{3}$ or $x_{2}\sim x_{4}$, then $x_{1}$ would have a 3-triangle. So suppose $x_{1}\not\sim x_{3}$ and $x_{2}\not\sim x_{4}$. By Proposition 47, this means that $x_{2},x_{4}$ are the only shared neighbors of $x_{1},x_{3}$. If $\deg(x_{3})=4$, then we would have $\operatorname{shad}(N[x_{3}])\leq 0$ (due to 1-vertex $x_{1}$), contradicting that $G$ is simplified. So, suppose that $\deg(x_{3})=5$. Since $x_{2},x_{4}$ are neighbors of 5-vertex $x_{3}$ and 3-vertex $x_{1}$, they must have degree 4 (as 5-vertices cannot be neighbors of 3-vertices). At this point, observe that $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$) and $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 1$ (due to 1-vertex $x_{1}$), contradicting that $G$ is in 5BP.
For the second result, suppose $\deg(x_{1})=4$ and $x_{1}\sim x_{3}$ and $\deg(x_{2})\geq\deg(x_{4})$; we have already shown that $\deg(x_{4})\geq 4$. We have $x_{2}\not\sim x_{4}$, as otherwise $x_{1}$ has a funnel (due to clique $x_{2},x_{3},x_{4}$). If $\deg(x_{2})=5$, then $R(x_{2})\geq 1$ (due to 3-triangle $x_{1},x_{3},x_{4}$). Otherwise, if $\deg(x_{2})=\deg(x_{4})=4$, then $R(N(x_{1})\cap N(x_{3}))\geq 1$ (due to subquadratic vertex $x_{2}$) and $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$); this contradicts that $G$ is in 5BP.
For the final result, suppose that $\deg(x_{1})=\deg(x_{2})=\deg(x_{3})=4$. We have already shown then that $x_{1}\not\sim x_{3}$, and Proposition 47 gives $\operatorname{codeg}(x_{1},x_{3})\leq 2$. So $G-\{x_{2},x_{4}\}$ has two non-adjacent subquadratic vertices $x_{1},x_{3}$. Thus $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 2$ by Observation 16, and $R(x_{1},x_{3})\geq 1$ due to subquadratic vertex $x_{2}$. This contradicts that $G$ is in 5BP. ∎
@@
Suppose $G$ is in 5BP. If a 4-vertex $u$ has a 5-neighbor, then the neighbors of $u$ can be enumerated as $v_{1},v_{2},v_{3},v_{4}$, where $\deg(v_{1})=5$ and $\{v_{3},v_{4}\}\not\sim\{v_{1},v_{2},v_{3},v_{4}\}$. Moreover, after Line 4, we may further suppose that $v_{3},v_{4}$ are subquartic.
@@
Let $v_{1}$ be a 5-neighbor of $u$. By Proposition 48, we must have $\operatorname{codeg}(u,v_{1})\leq 1$. Also, if there are any edges among the other neighbors of $u$, then $R(v_{1})\geq 1$ due to the resulting 3-triangle on $u$. Finally, if either $v_{3}$ or $v_{4}$ has degree 5, then $u$ would have two non-adjacent 5-neighbors and we would branch at Line 4. ∎
@@
Suppose $x\sim u$ and $u,x$ share neighbors $y_{1},y_{2}$; let $v_{1},v_{2}$ be the two remaining neighbors of $u$, where $\deg(v_{1})=5$. Since we did not branch on $v_{1},v_{2}$ at Line 3, we must have $\deg(v_{2})=4,v_{1}\sim v_{2}$ and $\deg(y_{3})=\deg(y_{2})=\deg(x)=5$. By Proposition 49, we have $\{y_{1},y_{2},x\}\not\sim v_{2}$. Then $y_{1}\not\sim v_{1}$, as otherwise $G-\{y_{2},x\}$ would have a kite, where $\deg_{G-y_{2}}(v_{2})=\deg_{G}(v_{2})=4$, and we would branch at Line 3. So $R(v_{1},y_{1})\geq 1$ (due to the 3-triangle $u,x,y_{2}$) and $R(v_{1},N[y_{1}])\geq 1$ (due to 2-vertex $v_{2}$), and we would branch on non-adjacent 5-vertices $v_{1},y_{1}$ at Line 2. ∎
@@
Let $A_{51},A_{5+}$ be the set of 5-vertices with exactly one 4-neighbor or more than one 4-neighbor, respectively, and let $B$ be the set of 4-vertices with exactly one 5-neighbor. Note that $A_{51},A_{5+}$ partition $V_{5}$ at this point; for, if a 5-vertex $u$ has five independent 5-neighbors, we would branch on it at Line 6; if it has a triangle in its neighborhood, we would branch on $u$ at Line 5.
We claim that, for any vertex $u\in A_{51}$ with 4-neighbor $x$, we have $x\in B$. For, suppose that $x$ has a second 5-neighbor $v$; we must have $u\sim v$ else we would branch on $x$ at Line 4. Then $u$ has three other 5-neighbors $v_{3},v_{4},v_{5}$; we would branch on $u$ at Line 5.
By counting edges between $B$ and $A_{51}$, we thus see that $|B|\geq|A_{51}|$. Next let us count the edges between $V_{4}$ and $V_{5}$. Clearly there are least $|A_{51}|+2|A_{5+}|$ such edges. Furthermore, each vertex $v\in B$ has exactly one 5-neighbor, and by Proposition 49 any 4-vertex $v$ has at most two 5-neighbors. So $|B|+2(|V_{4}|-|B|)\leq|A_{51}|+2|A_{5+}|$. Combining this inequality with $|B|\geq|A_{51}|$, we conclude that $|V_{4}|=|B|+(|V_{4}|-|B|)\geq|A_{51}|+|A_{5+}|=|V_{5}|.$
Observation 52.
@@
If $\operatorname{shad}(N[x])\leq 1$, then since $G$ is in 5BP we would have $R(N[x])\geq 1$ and we would branch on $x$ at Line 2. So $\deg_{G-N[x]}(y)\geq 3$, i.e. $\operatorname{codeg}(x,y)\leq\deg(y)-3$. We have already seen that $x$ has no 3-neighbor since $G$ is in 5BP; this also shows it cannot have any 3-vertex at distance two. ∎
@@
Let $G^{\prime}=G-N[x,y]$. We first claim that $\operatorname{mindeg}(G^{\prime})\geq 2$. For, consider a vertex $t\in G^{\prime}$. If $\deg_{G}(t)=3$, then by Observation 52 we have $\operatorname{codeg}(x,t)=0$ and by Proposition 48 we have $\operatorname{codeg}(y,t)\leq 1$. So $\deg_{G^{\prime}}(t)\geq 2$. Likewise, if $\deg_{G}(t)=5$, then $\operatorname{codeg}(y,t)\leq 1$ and $\operatorname{codeg}(x,t)\leq 2$. So again $\deg_{G^{\prime}}(t)\geq 2$.
Next suppose $\deg_{G}(t)=4$; by Proposition 47 and Observation 52 we have $\operatorname{codeg}(y,t)\leq 2$ and $\operatorname{codeg}(x,t)\leq 1$. To get $\deg_{G^{\prime}}(t)=1$, vertex $t$ must share two neighbors $z_{1},z_{2}$ with $y$ and share a distinct neighbor $z_{3}$ with $x$. By Proposition 48 we must have $\deg(z_{1})=\deg(z_{2})=5$. If $z_{1}\not\sim z_{2}$, then $R(z_{1},z_{2})\geq 1$ (due to subquadratic vertex $y$), and we would branch on $z_{1},z_{2}$ at Line 9. If $z_{1}\sim z_{2}$, then in the graph $G-N[x]$, vertex $t$ loses its neighbor $z_{3}$, and so it has a 3-triangle $z_{1},z_{2}$; in particular, $R(N[x])\geq 1$ and we would branch on $x$ at Line 8.
Thus, in all cases, $\deg_{G^{\prime}}(t)\geq 2$. Now, suppose $G^{\prime}$ has an indset $I$ with $\operatorname{surp}_{G^{\prime}}(I)\leq 0$. Since $\operatorname{mindeg}(G^{\prime})\geq 2$, necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 0+\deg_{G-N[x]}(y)-1=2$. So $G-N[x]$ would be 3-vul and we would branch on $x$ at Line 8. ∎
Observation 54.
After Line 10 of algorithm Branch5-2, any 4-vertex has at most one 5-neighbor and is not part of any 4-cycle.
@@
Consider a 4-vertex $v$. If $v$ has two 5-neighbors we would branch at Line 6 or Line 10, depending if they were adjacent or not. Now consider a four-cycle $v,x_{1},x_{2},x_{3}$. By Observation 52, we have $\deg(x_{2})\leq 4$. Thus, by Proposition 48, both $x_{1}$ and $x_{3}$ have degree 5; that is, $v$ has two 5-neighbors, which is what we have showed is impossible.
@@ 9.2 Basic branching rules
Line 1: We can first simplify $G$. If a vertex $u$ has degree at least 6, we apply Section 5. If a 5-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$ we apply Lemma 26. Otherwise if a 5-vertex $u$ has $R(u)\geq 1$, then we split on $u$ giving branch-seq $[(0.5,2),(2,5)]$. If a 4-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$, then we apply Proposition 25. If vertices $u,v$ have $R(u,v)+R(N(u)\cap N(v)))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
@@
Line 2: We split on $u,v$, obtaining subproblems $G_{1}=G-u,G_{2}=G-u-N[v],G_{3}=G-\{u,v\}$. Since $G$ is in 5BP, they have drop $(2,5),(2.5,6),(1,2)$ directly. After simplification, $G_{2}$ has drop $(2.5,7)$ and $G_{3}$ has drop $(1,3)$. So we need:
@@
Line 3: We split on $u,v$ and apply (P3) to $G-\{u,v\}$. Since $G$ is in 5BP, subproblems $G-N[u]$ and $G-u-N[v]$ both have drops $(2,5)$. Also, $\operatorname{shad}(u,v)\geq\operatorname{shad}(u)-1\geq 1$ so by Lemma 9 the subproblem $G-\{u,v\}$ has drop $(1.5,4)$ after simplification. We need:
@@
Line 4: Consider a 4-vertex $u$ with non-adjacent 5-neighbors $y_{1},y_{2}$, and let $x_{1},x_{2}$ be the other two neighbors of $u$ where $\deg(x_{1})\leq\deg(x_{2})$. Our plan is to split on $y_{1},y_{2}$, generating subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}$. Since $G$ is in 5BP, the subproblems have drops $(2,5),(2.5,6),(1,2)$ directly.
By Proposition 49, the neighbors of $u$ must be independent. If $\deg(x_{1})\leq 3$, then $R(y_{1},y_{2})\geq 1$ (due to 2-vertex $u$) and $R(y_{1},N[y_{2}])\geq 1$ (due to 2-vertex $x_{1}$); in this case we would branch on $y_{1},y_{2}$ at Line 2. So we may assume that $\deg(x_{1})\geq 4$. We can apply Proposition 30 to 2-vertex $u$ in $G_{3}$ with $r=\deg(x_{1})+\deg(x_{2})-\operatorname{codeg}(x_{1},x_{2})-1$. If $\deg(x_{1})=\deg(x_{2})=4$, then Proposition 48 gives $\operatorname{codeg}(x_{1},x_{2})=1$ (they share 4-neighbor $u$). If $\deg(x_{2})=5$, then Proposition 47 gives $\operatorname{codeg}(x_{1},x_{2})\leq 2$. In either case $r\geq 6$ and so we need:
@@
Line 5: Vertices $v_{3},v_{4},v_{5}$ cannot form a clique (else $R(v_{1})\geq 1$ due to the funnel $u$), so assume without loss of generality $v_{3}\not\sim v_{4}$. Then we split on $v_{3},v_{4}$, generating subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}$, which have drops $(2,5),(2.5,6),(1,2)$ directly.
In $G_{3}$, we also apply Proposition 30 to vertices $v_{1},v_{5}$ for the 3-triangle $u,v_{1},v_{2}$. By Proposition 50, at this point $\operatorname{codeg}(v_{1},u)\leq 1$ and $\operatorname{codeg}(v_{5},u)\leq 1$. So $v_{1}\not\sim v_{5}$ and $\deg_{G_{3}}(v_{1})=5,\deg_{G_{3}}(v_{5})\geq 4$ and, by Proposition 47, we have $\operatorname{codeg}(v_{1},v_{5})\leq 2$. So $\deg_{G_{3}}(v_{1})+\deg_{G_{3}}(v_{5})-\operatorname{codeg}_{G_{3}}(v_{1},v_{5})-1\geq 5+4-2-1=6$. Overall we need
@@
Line 8: Since $G$ is in 5BP, subproblem $G-N[u]$ has drop $(2,5)$ directly, and subproblem $G-u$ has drop $(0.5,1)$. We apply Proposition 31 to $G-N[u]$, so we need:
@@
Line 9: We get subproblems $G-N[u],G-u-N[v],G-\{u,v\}$. Since $G$ is in 5BP, these have drops $(2,5),(2.5,6),(1,2)$ directly, and the latter has drop $(1,3)$ after simplification. We need:

### v7

@@ 7.1 Regularizing the graph and structural properties
In Line 1, we will apply a series of branching rules to ensure the following conditions:
@@
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $S(u)=0$.
@@
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Note that, when these conditions hold, then no 5-vertex $u$ can have subcubic neighbors (else we would have $S(u)\geq 1$). Unlike in the degree-4 case, we cannot ensure that $G$ has no four-cycles. However, we can show that the four-cycles are heavily restricted.
@@
After Line 1, any pair of non-adjacent vertices $x,y$ have $\operatorname{codeg}(x,y)\leq 2$.
@@
Suppose $\deg(x)\geq\deg(y)$ and $\operatorname{codeg}(x,y)\geq 3$. If $\deg(x)=3$, then $\{x,y\}$ would be an indset with surplus one, contradicting that $G$ is simplified. If $\deg(y)=4$, then $\operatorname{shad}(N[y])\leq 0$ (due to vertex $x$), and we would branch at Line 1. So we suppose $\deg(x)=\deg(y)=5$. Again, if $\operatorname{codeg}(x,y)\geq 4$ then $\operatorname{shad}(N[y])\leq 0$ due to vertex $x$, so we suppose that $\operatorname{codeg}(x,y)=3$ exactly. Then $S(N(x)\cap N(y))\geq 2$ due to non-adjacent 2-vertices $x,y$. So $S(x,y)+S(N(x)\cap N(y))+\operatorname{codeg}(x,y)\geq 5$ and we would have branched at Line 1. ∎
@@
After Line 1, any 4-cycle $x_{1},x_{2},x_{3},x_{4}$ satisfies the following properties:
@@
For the first result, suppose that $\deg(x_{1})=3$. If $x_{1}\sim x_{3}$ or $x_{2}\sim x_{4}$, then $x_{1}$ would have a 3-triangle. So suppose $x_{1}\not\sim x_{3}$ and $x_{2}\not\sim x_{4}$. By Proposition 48, this means that $x_{2},x_{4}$ are the only shared neighbors of $x_{1},x_{3}$. If $\deg(x_{3})=4$, then $\operatorname{shad}(N[x_{3}])\leq 0$ (due to 1-vertex $x_{1}$), contradicting that $G$ is simplified. So, suppose that $\deg(x_{3})=5$. Since $x_{2},x_{4}$ are neighbors of 5-vertex $x_{3}$ and 3-vertex $x_{1}$, they must have degree 4 (as 5-vertices cannot be neighbors of 3-vertices). At this point, observe that $S(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$) and $S(N(x_{1})\cap N(x_{3}))=S(x_{2},x_{4})\geq 1$ (due to 1-vertex $x_{1}$); we would branch at Line 1.
For the second result, suppose without loss of generality that $\deg(x_{1})=4$ and $x_{1}\sim x_{3}$ and $\deg(x_{2})\geq\deg(x_{4})$; we have already shown that $\deg(x_{4})\geq 4$. We have $x_{2}\not\sim x_{4}$, as otherwise $x_{1}$ has a funnel (due to clique $x_{2},x_{3},x_{4}$). If $\deg(x_{2})=5$, then $S(x_{2})\geq 1$ (due to 3-triangle $x_{1},x_{3},x_{4}$). Otherwise, if $\deg(x_{2})=\deg(x_{4})=4$, then $S(N(x_{1})\cap N(x_{3}))\geq 1$ (due to subquadratic vertex $x_{2}$) and $S(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$); we would branch at Line 1.
For the final result, suppose that $\deg(x_{1})=\deg(x_{2})=\deg(x_{3})=4$. We have already shown then that $x_{1}\not\sim x_{3}$, and Proposition 48 gives $\operatorname{codeg}(x_{1},x_{3})\leq 2$. So $G-\{x_{2},x_{4}\}$ has two non-adjacent subquadratic vertices $x_{1},x_{3}$. Thus $S(N(x_{1})\cap N(x_{3}))=S(x_{2},x_{4})\geq 2$ by Observation 12, and $S(x_{1},x_{3})\geq 1$ due to subquadratic vertex $x_{2}$. We would branch at Line 1. ∎
@@
After Line 1, the neighbors of any 4-vertex $u$ can be enumerated as $v_{1},\dots,v_{4}$ to satisfy one of the following three conditions:
  • (I)
    $v_{1},\dots,v_{4}$ are all subquartic.
  • (II)
    $\deg(v_{1})=5$ and $v_{3},v_{4}$ are subquartic, and $G[N(u)]$ has exactly one edge $(v_{1},v_{2})$.
  • (III)
    $\deg(v_{1})=5$, and all the vertices $v_{1},\dots,v_{4}$ are independent.
Moreover, after Line 4, in case (III) we may further suppose that $v_{2},v_{3},v_{4}$ are subquartic.
@@
Suppose $u$ has a 5-neighbor $v_{1}$. By Proposition 49, we must have $\operatorname{codeg}(u,v_{1})\leq 1$. Also, if there are any edges among the other neighbors of $u$, then $S(v_{1})\geq 1$ due to the resulting 3-triangle on $u$. In Case II, if say $\deg(v_{3})=5$, then we would have $S(v_{3})\geq 1$ (due to the 3-triangle on $u,v_{1},v_{2}$), and we would have applied the Line-1 branching rules. In Case III, if say $\deg(v_{2})=5$, then $u$ would have two non-adjacent 5-neighbors and we would branch at Line 4. ∎
@@
Suppose $x\sim u$ and $u,x$ share neighbors $y_{1},y_{2}$; let $v_{1},v_{2}$ be the two remaining neighbors of $u$, where $\deg(v_{1})=5$. Note that $G-\{v_{1},v_{2}\}$ contains a kite $u,x,y_{1},y_{2}$. Since we did not branch on $v_{1},v_{2}$ at Line 3, we must have $\deg(v_{2})=4,v_{1}\sim v_{2}$. Since $u$ has at least four 5-neighbors, this implies that $\deg(y_{1})=\deg(y_{2})=\deg(x)=5$.
The 4-vertex $v_{2}$ must be in Case II of Proposition 50. So $\{y_{1},y_{2},x\}\not\sim v_{2}$. Then $y_{1}\not\sim v_{1}$, as otherwise $G-\{y_{2},v_{2}\}$ would have a kite, where $\deg_{G-y_{2}}(v_{2})=4$, and we would branch at Line 3. So $S(v_{1},y_{1})\geq 1$ (due to the 3-triangle $u,x,y_{2}$) and $S(v_{1},N[y_{1}])\geq 1$ (due to 2-vertex $v_{2}$), and we would branch on non-adjacent 5-vertices $v_{1},y_{1}$ at Line 2. ∎
@@
Let $A_{51},A_{5+}$ be the set of 5-vertices with exactly one 4-neighbor or more than one 4-neighbor respectively, and let $B$ be the set of 4-vertices with exactly one 5-neighbor. Note that $A_{51},A_{5+}$ partition $V_{5}$ at this point; for, if a 5-vertex $u$ has five independent 5-neighbors, we would branch on it at Line 6; if it has a triangle in its neighborhood, we would branch on $u$ at Line 5.
We claim that, for any vertex in $A_{51}$ its sole 4-neighbor must be in $B$, and hence $|B|\geq|A_{51}|$. For, suppose that $u\in A_{51}$ has a neighbor $x\in B$ and $x$ has a second 5-neighbor $v$; if $u\not\sim v$ we would branch on $x$ at Line 4 and if $u\sim v$ we would branch on $u$ at Line 5.
There are at least $|A_{51}|+2|A_{5+}|$ edges between $V_{4}$ and $V_{5}$. Conversely, each vertex $v\in B$ has exactly one 5-neighbor, and by Proposition 50 any 4-vertex $v$ has at most two 5-neighbors. So $|B|+2(|V_{4}|-|B|)\leq|A_{51}|+2|A_{5+}|$. Since $|B|\geq|A_{51}|$, we conclude that $|V_{4}|=|B|+(|V_{4}|-|B|)\geq|A_{51}|+|A_{5+}|=|V_{5}|.$
Observation 53.
@@
If $\operatorname{shad}(N[x])\leq 1$, then after Line 1 we would have $S(N[x])\geq 1$ and we would branch on $x$ at Line 8. So $\deg_{G-N[x]}(y)\geq 3$, i.e. $\operatorname{codeg}(x,y)\leq\deg(y)-3$. We have already seen that $x$ has no 3-neighbor; this also shows it cannot have any 3-vertex at distance two. ∎
@@
Let $G^{\prime}=G-N[x,y]$. We first claim that $\operatorname{mindeg}(G^{\prime})\geq 2$. For, consider a vertex $t\in G^{\prime}$. If $\deg_{G}(t)=3$, then by Observation 53 we have $\operatorname{codeg}(x,t)=0$ and by Proposition 49 we have $\operatorname{codeg}(y,t)\leq 1$. So $\deg_{G^{\prime}}(t)\geq 2$. Likewise, if $\deg_{G}(t)=5$, then $\operatorname{codeg}(y,t)\leq 1$ and $\operatorname{codeg}(x,t)\leq 2$. So again $\deg_{G^{\prime}}(t)\geq 2$.
Next suppose $\deg_{G}(t)=4$; by Proposition 48 and Observation 53 we have $\operatorname{codeg}(y,t)\leq 2$ and $\operatorname{codeg}(x,t)\leq 1$. To get $\deg_{G^{\prime}}(t)=1$, vertex $t$ must share two neighbors $z_{1},z_{2}$ with $y$ and share a distinct neighbor $z_{3}$ with $x$. By Proposition 49 we must have $\deg(z_{1})=\deg(z_{2})=5$. If $z_{1}\not\sim z_{2}$, then $S(z_{1},z_{2})\geq 1$ (due to subquadratic vertex $y$), and we would branch on $z_{1},z_{2}$ at Line 9. If $z_{1}\sim z_{2}$, then in the graph GN[x]G-N[x], vertex $t$ loses its neighbor $z_{3}$, and so it has a 3-triangle $z_{1},z_{2}$; in particular, $S(N[x])\geq 1$ and we would branch on $x$ at Line 8.
Thus, in all cases, $\deg_{G^{\prime}}(t)\geq 2$. Now, suppose $G^{\prime}$ has an indset $I$ with $\operatorname{surp}_{G^{\prime}}(I)\leq 0$. Since $\operatorname{mindeg}(G^{\prime})\geq 2$, necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 0+\deg_{G-N[x]}(y)-1=2$. So GN[x]G-N[x] would be 3-vul and we would branch on $x$ at Line 8. ∎
Observation 55.
After Line 10 of algorithm Branch5-2, any 4-vertex $v$ has at most one 5-neighbor and is not part of any 4-cycle.
@@
If $v$ has two 5-neighbors we would branch at Line 4 or Line 10, depending if they were adjacent or not. Now consider a four-cycle $v,x_{1},x_{2},x_{3}$. By Observation 53, we have $\deg(x_{2})\leq 4$. Thus, by Proposition 49, we have $\deg(x_{1})=\deg(x_{3})=5$; that is, $v$ has two 5-neighbors, which is what we have showed is impossible. ∎
@@ 7.2 Basic branching rules
Line 1: We first simplify $G$. If a vertex has degree at least 6, we apply Figure 6. If a 5-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$ we apply Lemma 26. Otherwise if a 5-vertex $u$ has $S(u)\geq 1$, then we split on $u$ giving branch-seq $[(0.5,2),(2,5)]$. If a 4-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$, then we apply Proposition 25. If vertices $u,v$ have $S(u,v)+S(N(u)\cap N(v)))+\operatorname{codeg}(u,v)\geq 5$, we apply (B-Shared).
@@
Line 2: We split on $u,v$, obtaining subproblems $G_{1}=G-u,G_{2}=G-u-N[v],G_{3}=G-\{u,v\}$. They have drop $(2,5),(2.5,6),(1,2)$ directly. After simplification, $G_{2}$ has drop $(2.5,7)$ and $G_{3}$ has drop $(1,3)$. So we need:
@@
Line 3: We split on $u,v$ and apply (P3) to $G-\{u,v\}$. Since we are after Line 1, subproblems GN[u]G-N[u] and GuN[v]G-u-N[v] both have drops $(2,5)$. Also, $\operatorname{shad}(u,v)\geq\operatorname{shad}(u)-1\geq 1$ so by Lemma 14 the subproblem $G-\{u,v\}$ has drop $(1.5,4)$ after simplification. We need:
@@
Line 4: Consider a 4-vertex $u$ with non-adjacent 5-neighbors $y_{1},y_{2}$. This must be Case III of Proposition 50, where all neighbors of $u$ are independent; let $x_{1},x_{2}$ be the other two neighbors of $u$, sorted so that $\deg(x_{1})\leq\deg(x_{2})$. We split on $y_{1},y_{2}$, generating subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}$. Since we are after Line 1, the subproblems have drops $(2,5),(2.5,6),(1,2)$ directly.
If $\deg(x_{1})=3$, then $S(y_{1},y_{2})\geq 1$ (due to 2-vertex $u$) and $S(y_{1},N[y_{2}])\geq 1$ (due to 2-vertex $x_{1}$); in this case we would branch on $y_{1},y_{2}$ at Line 2. So suppose that $\deg(x_{1})\geq 4$. We can apply Proposition 30 to 2-vertex $u$ in $G_{3}$ with $r=\deg(x_{1})+\deg(x_{2})-\operatorname{codeg}(x_{1},x_{2})-1$. If $\deg(x_{1})=\deg(x_{2})=4$, then Proposition 49 gives $\operatorname{codeg}(x_{1},x_{2})=1$ (they share 4-neighbor $u$). If $\deg(x_{2})=5$, then Proposition 48 gives $\operatorname{codeg}(x_{1},x_{2})\leq 2$. In either case $r\geq 6$ and so we need:
@@
Line 5: Vertices $v_{3},v_{4},v_{5}$ cannot form a clique (else $S(v_{1})\geq 1$ due to the funnel $u$), so assume without loss of generality $v_{3}\not\sim v_{4}$. Then we split on $v_{3},v_{4}$, generating subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}$, which have drops $(2,5),(2.5,6),(1,2)$ directly.
In $G_{3}$, we also apply Proposition 30 to vertices $v_{1},v_{5}$ for the 3-triangle $u,v_{1},v_{2}$. By Proposition 51, at this point $\operatorname{codeg}(v_{1},u)\leq 1$ and $\operatorname{codeg}(v_{5},u)\leq 1$. So $v_{1}\not\sim v_{5}$ and $\deg_{G_{3}}(v_{1})=5,\deg_{G_{3}}(v_{5})\geq 4$ and, by Proposition 48, we have $\operatorname{codeg}(v_{1},v_{5})\leq 2$. So $\deg_{G_{3}}(v_{1})+\deg_{G_{3}}(v_{5})-\operatorname{codeg}_{G_{3}}(v_{1},v_{5})-1\geq 6$. Overall we need
@@
Line 8: Since we are after Line 1, subproblem GN[u]G-N[u] has drop $(2,5)$ directly, and subproblem GuG-u has drop $(0.5,1)$. We apply Proposition 31 to GN[u]G-N[u], so we need:
@@
Line 9: We get subproblems $G-N[u],G-u-N[v],G-\{u,v\}$. Since we are after Line 1, these have drops $(2,5),(2.5,6),(1,2)$ directly, and the latter has drop $(1,3)$ after simplification. We need: